Note that yı(t) = Vt and yz(t) = t-1 are solutions of the linear homogeneous differential equation 2t’y + 3ty – y = 0. Use
Answer #1

(1) Given differential equation is 21y+3ty- y = 42 +41 Rewrite differential equation is (21+ D2 +34D-1)y= 0 This is CauchyTo find Particular Integral (P.1): Here yı = e. and y= e - And Wronski an of y1, y, is W = by yil (***)-(e){}) - Thus (PR)-